below are a rectangle and a right triangle.
The perimeter of the rectangle is 34 cm. The perimeter of the triangle is 30 cm. Find the values of m and n.
On the right triangle, the base is n and the side is m. The slope however is n+1. The rectangles base is n and the side is m.

Answer by edjones(7569)

2n+2m=34
.
n+m+n+1=30
2n+m=29
m=29-2n
2n+2(29-2n)=34 substitute 29-2n for m.
2n+58-4n=34
-2n=-24
n=12
m=5
.
Ed